COMMUTATIVE PARTIAL DIFFERENTIAL OPERATORS

ALEX KASMAN EMMA PREVIATO

Abstract. In one variable, there exists a satisfactory classification of commutative rings of differential operators. In several variables, even the simplest generalizations seem to be unknown and in this report we give examples and pose questions that may suggest a theory to be developed. In particular, we address the existence of a “spectral variety” generalizing the spectral curve of the one dimensional theory and the role of the differential resultant.